Probability:

if the omega - {A,B,C} is the sample space for an experiment then P({A}) = 1/3 (True or False?) explain.

if the omega - {A,B,C} is the sample space for an experiment, then P({A,B}) = 1 (True or False?) explain.

if a coin is tossed three times (and it is a fair coin) then the probability that H and T each come up at least once is 3/4 (True or False?) explain.
Answer #1

if the omega - {A,B,C} is the sample space for an experiment then P({A}) = 1/3

This is a false statement.

As there nowhere is written that all the three probabilities should be equal.

if the omega - {A,B,C} is the sample space for an experiment, then P({A,B}) = 1

This is a false statement.

As there nowhere is written that A and B always occur together

if a coin is tossed three times (and it is a fair coin) then the probability that H and T each come up at least once is 3/4

The statement is true

There are 8 possibilities in total and there are only two cases where only head {HHH} or only tails {TTT} occurs.

6 possibilities are there with atleast one head and tails.

probability = 6/8 = 3/4